# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 5.1. Left-shifts and right-shifts
The left-shift $(A^{\prime},B^{\prime})$ of a mixed-separation+ $(A,B)$ satisfies $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.∎
@@
Let $(A,B)$ be a mixed-$k$-separation+ of a graph $G$ with left-shift $(A^{\prime},B^{\prime})$.
@@
    $(A^{\prime},B^{\prime})$ is a $-\ell$-separation+ with $\ell\leqslant k$. We have $\ell=k$ if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
@@
(i). By definition, the left-shift $(A^{\prime},B^{\prime})$ is a mixed-separation+. The separator of $(A^{\prime},B^{\prime})$ is obtained from the separator of $(A,B)$ by replacing every vertex $v\in S(A,B)$ that has exactly one neighbour $u\in A\smallsetminus B$ with the edge $uv$ and by removing every vertex in $S(A,B)$ that has no neighbour in $A\smallsetminus B$. Hence the order of $(A^{\prime},B^{\prime})$ is at most that of $(A,B)$, with equality if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
(ii). Assume for a contradiction that there is a vertex $v\in S(A^{\prime},B^{\prime})$ with at most one neighbour in $A^{\prime}\smallsetminus B^{\prime}$. Since $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ by 5.2, the vertex $v$ also has at most one neighbour in $A\smallsetminus B$. But then, by the definition of left-shift, we have $v\notin A^{\prime}$, contradicting $v\in S(A^{\prime},B^{\prime})$. ∎
@@
(i) $\Rightarrow$ (ii) is immediate.
(ii) $\Rightarrow$ (i). Assume that (i) fails, so there are two edges $ab_{1},ab_{2}\in S(A,B)$ with $a\in A\smallsetminus B$ and both $b_{i}\in B\smallsetminus A$. Let $X$ be the vertex-set that consists of $a$ together with all vertices in $S(A,B)$ and all endvertices in $B$ of edges in $S(A,B)$ other than the $ab_{i}$. Since $|A\smallsetminus B|\geqslant 2$ by (ii), there is $\alpha\in A\smallsetminus B$ distinct from $a$. If some $b_{i}$ is not in $X$, we let $\beta:=b_{i}$. Otherwise both $b_{i}$ are in $X$, each has exactly two neighbours in $A\smallsetminus B$, and $S(A,B)$ contains only edges. Since $b_{1}$ has degree four, it has a neighbour $\beta$ in $B\smallsetminus A$ that is distinct from $b_{2}$. In either case, $X$ is a 3-separator that separates $\alpha$ from $\beta$, contradicting that $G$ is 4-connected. ∎
@@
(i) $\Rightarrow$ (ii) is immediate.
(ii) $\Rightarrow$ (iii). Assume $|A\smallsetminus B|\leqslant 1$. Since every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, we get that $S(A,B)$ is an edge-cut. Since $S(A,B)$ is nonempty, $|A|=1$ follows.
(iii) $\Rightarrow$ (i). This follows from Lemma 5.4 (ii) $\Rightarrow$ (i). ∎
@@
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$ with $A\smallsetminus B\neq\emptyset$ and such that no two edges in $S(A,B)$ share an end in $B\smallsetminus A$. Let $(A^{\prime},B^{\prime})$ denote the left-shift of $(A,B)$. Then the following assertions are equivalent:
@@
By Lemma 5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
(i) $\Rightarrow$ (ii). Suppose $|A\smallsetminus B|\leqslant 1$. So $|A\smallsetminus B|=1$ since $A\smallsetminus B\neq\emptyset$. Then $|A^{\prime}\smallsetminus B^{\prime}|=1$ since $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$. Applying Lemma 5.5 to $(A^{\prime},B^{\prime})$ yields that $(A^{\prime},B^{\prime})$ is an atomic edge-cut.
(ii) $\Rightarrow$ (i). From $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$ we get $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$. So no two edges in $S(A^{\prime},B^{\prime})$ share an end in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma 5.5 applied to $(A^{\prime},B^{\prime})$. So suppose for a contradiction that there are two distinct edges $a_{1}b,a_{2}b$ in $S(A^{\prime},B^{\prime})$ with $a_{i}\in A^{\prime}\smallsetminus B^{\prime}$ and $b\in B^{\prime}\smallsetminus A^{\prime}$. Then $a_{i}\in A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $b\in B\cap A$. Hence $b$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, contradicting that $b$ was replaced with edges when taking the left-shift. ∎
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ be a mixed-4-separation of $G$. Let $(C,D)$ be a 4-separation of $G$ with right-shift $(C^{\prime},D^{\prime})$. Assume $A\subseteq C$. Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
@@
We obtain $A\subseteq C^{\prime}$ from $A\subseteq C=C^{\prime}$. Let $X$ consist of all endvertices in $A$ of edges in $S(A,B)$. Then $N(B\smallsetminus A)=(A\cap B)\cup X$. From $A\subseteq C$ we obtain $D\smallsetminus C\subseteq B\smallsetminus A$. Hence $D\subseteq(B\smallsetminus A)\cup N(B\smallsetminus A)=B\cup X$. Recall that $D^{\prime}\subseteq D$. Let $v\in D^{\prime}$. It suffices to show that $v\notin X$. If $v\in X$, then $v$ has exactly one neighbour in $B\smallsetminus A$. Since $B\smallsetminus A\supseteq D\smallsetminus C$, the definition of right-shift yields $v\notin D^{\prime}$, a contradiction. So $v\notin X$. ∎
5.2. Left-right-shifts and right-left-shifts
@@
Let $(A,B)$ be a mixed-separation+. The left-right-shift of $(A,B)$ is obtained from $(A,B)$ by first taking the left-shift and then the right-shift. The right-left-shift of $(A,B)$ is obtained from $(A,B)$ by first taking the right-shift and then the left-shift.
@@
Let $(A^{\prime},B^{\prime})$ be the left-right-shift of a mixed-4-separation+ $(A,B)$ in a graph $G$. Then $A^{\prime}\smallsetminus B^{\prime}\supseteq A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.
@@
We perform the two shifts one after the other, applying 5.2 after each shift. ∎
@@
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$. The left-right-shift $(A^{\prime\prime},B^{\prime\prime})$ of $(A,B)$ is a tetra-separation if and only if the following three conditions are satisfied:
@@
Let $(A^{\prime},B^{\prime})$ denote the left-shift of $(A,B)$.
@@
Backward implication. Assume (i)(iii). By (iii) and Lemma 5.4 (applied to $(B,A)$ instead of $(A,B)$), no two edges in $S(A,B)$ share an endvertex in $B\smallsetminus A$. By (i) and Lemma 5.6, $(A^{\prime},B^{\prime})$ satisfies the matching-condition. By Lemma 5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$. By (i) and since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$. Hence $(A^{\prime},B^{\prime})$ has order four by Lemma 5.3. We further have $|A^{\prime}\smallsetminus B^{\prime}|=|A\smallsetminus B|\geqslant 2$ by 5.9 and we have $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$ by (ii).
@@
By Lemma 5.6 (applied for the right-shift), $(A^{\prime\prime},B^{\prime\prime})$ satisfies the matching-condition. By Lemma 5.3 (applied for the right-shift), every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $B^{\prime\prime}\smallsetminus A^{\prime\prime}$. Every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $A^{\prime\prime}\smallsetminus B^{\prime\prime}$ since $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ by 5.9 (applied for the right-shift). Hence $(A^{\prime\prime},B^{\prime\prime})$ satisfies the degree-condition. Since $B^{\prime}\smallsetminus A^{\prime}$ is nonempty and $G$ is 4-connected, every vertex in $S(A^{\prime},B^{\prime})$ has a neighbour in $B^{\prime}\smallsetminus A^{\prime}$, so $(A^{\prime\prime},B^{\prime\prime})$ has order four by Lemma 5.3 (applied for the right-shift). Finally, $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by 5.9, so $(A^{\prime\prime},B^{\prime\prime})$ is proper.
@@
Forward implication. Suppose first that (iii) fails. Then, by Lemma 5.4 for $(B,A)$, there are two edges $ab_{1},ab_{2}$ in $S(A,B)$ with $a\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$. Then $a\in A^{\prime\prime}\smallsetminus B^{\prime\prime}$ and $b_{i}\in B^{\prime\prime}\smallsetminus A^{\prime\prime}$ by 5.9. Hence $(A^{\prime\prime},B^{\prime\prime})$ fails the matching-condition.
@@
Suppose next that (iii) holds, but (i) or (ii) fails. By (iii) and Lemma 5.4 for $(B,A)$, no two edges in $S(A^{\prime},B^{\prime})$ share an endvertex in $B^{\prime}\smallsetminus A^{\prime}$. If (i) fails, we have $|A\smallsetminus B|\leqslant 1$, so Lemma 5.6 gives that $(A^{\prime},B^{\prime})$ violates the matching-condition, and hence $(A^{\prime\prime},B^{\prime\prime})$ also violates the matching-condition. Next, assume that (i) holds, but (ii) fails. As $|A\smallsetminus B|\geqslant 2$, Lemma 5.6 gives that $(A^{\prime},B^{\prime})$ satisfies the matching-condition. The failure of (ii) yields $|B^{\prime}\smallsetminus A^{\prime}|\leqslant 1$. Hence applying Lemma 5.6 to $(B^{\prime},A^{\prime})$ (instead of $(A,B)$) yields that $(A^{\prime\prime},B^{\prime\prime})$ violates the matching-condition. ∎
Proof of Lemma 2.3..
(i). If $G$ is not quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Hence the right-left-shift of $(A,B)$ is a tetra-separation of $G$ by 5.10.
@@
(ii). Let $(A,B)$ be a tetra-separation of $G$. Let $X:=A\smallsetminus B$ and $Y:=B\smallsetminus A$. Let $F$ denote the set of edges in $S(A,B)$. We have
$8\leqslant|G|=|X|+|Y|+4-|F|$
@@
Both $X$ and $Y$ have size $\geqslant 2$ by Lemma 4.2 since $(A,B)$ is a tetra-separation. Hence both terms $|X|-2$ and $|Y|-2$ are $\geqslant 0$. We may therefore bipartition $F$ into classes $F_{X}$ and $F_{Y}$ of $\leqslant|X|-2$ and $\leqslant|Y|-2$ edges, respectively. We obtain $X^{\prime}$ from $X$ by adding for each edge in $F_{X}$ its endvertex in $A$ to $X$, and we obtain $Y^{\prime}$ similarly. Then $(X^{\prime},Y^{\prime})$ is a 4-separation of $G$. We have $|X^{\prime}\smallsetminus Y^{\prime}|\geqslant|X|-|F_{X}|\geqslant 2$, and similarly $|Y^{\prime}\smallsetminus X^{\prime}|\geqslant 2$. Thus $G$ is not quasi-5-connected. ∎

### v2

@@ 5.1. Left-reductions and right-reductions
The left-reduction $(A^{\prime},B^{\prime})$ of a mixed-separation+ $(A,B)$ satisfies $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.∎
@@
Let $(A,B)$ be a mixed-$k$-separation+ of a graph $G$ with left-reduction $(A^{\prime},B^{\prime})$.
@@
    $(A^{\prime},B^{\prime})$ is a mixed-$\ell$-separation+ with $\ell\leqslant k$. We have $\ell=k$ if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
@@
˜(i). By definition, the left-reduction $(A^{\prime},B^{\prime})$ is a mixed-separation+. The separator of $(A^{\prime},B^{\prime})$ is obtained from the separator of $(A,B)$ by replacing every vertex $v\in S(A,B)$ that has exactly one neighbour $u\in A\smallsetminus B$ with the edge $uv$ and by removing every vertex in $S(A,B)$ that has no neighbour in $A\smallsetminus B$. Hence the order of $(A^{\prime},B^{\prime})$ is at most that of $(A,B)$, with equality if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
˜(ii). Assume for a contradiction that there is a vertex $v\in S(A^{\prime},B^{\prime})$ with at most one neighbour in $A^{\prime}\smallsetminus B^{\prime}$. Since $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ by ˜5.2, the vertex $v$ also has at most one neighbour in $A\smallsetminus B$. But then, by the definition of left-reduction, we have $v\notin A^{\prime}$, contradicting $v\in S(A^{\prime},B^{\prime})$. ∎
@@
˜(i) $\Rightarrow$ ˜(ii) is immediate.
˜(ii) $\Rightarrow$ ˜(i). Assume that ˜(i) fails, so there are two edges $ab_{1},ab_{2}\in S(A,B)$ with $a\in A\smallsetminus B$ and both $b_{i}\in B\smallsetminus A$. Let $X$ be the vertex-set that consists of $a$ together with all vertices in $S(A,B)$ and all endvertices in $B$ of edges in $S(A,B)$ other than the $ab_{i}$. Since $|A\smallsetminus B|\geqslant 2$ by ˜(ii), there is $\alpha\in A\smallsetminus B$ distinct from $a$. If some $b_{i}$ is not in $X$, we let $\beta:=b_{i}$. Otherwise both $b_{i}$ are in $X$, each has exactly two neighbours in $A\smallsetminus B$, and $S(A,B)$ contains only edges. Since $b_{1}$ has degree four, it has a neighbour $\beta$ in $B\smallsetminus A$ that is distinct from $b_{2}$. In either case, $X$ is a 3-separator that separates $\alpha$ from $\beta$, contradicting that $G$ is 4-connected. ∎
@@
˜(i) $\Rightarrow$ ˜(ii) is immediate.
˜(ii) $\Rightarrow$ ˜(iii). Assume $|A\smallsetminus B|\leqslant 1$. Since every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, we get that $S(A,B)$ is an edge-cut. Since $S(A,B)$ is nonempty, $|A|=1$ follows.
˜(iii) $\Rightarrow$ ˜(i). This follows from Lemma˜5.4 ˜(ii) $\Rightarrow$ ˜(i). ∎
@@
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$ with $A\smallsetminus B\neq\emptyset$ and such that no two edges in $S(A,B)$ share an end in $B\smallsetminus A$. Let $(A^{\prime},B^{\prime})$ denote the left-reduction of $(A,B)$. Then the following assertions are equivalent:
@@
By Lemma˜5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
˜(i) $\Rightarrow$ ˜(ii). Suppose $|A\smallsetminus B|\leqslant 1$. So $|A\smallsetminus B|=1$ since $A\smallsetminus B\neq\emptyset$. Then $|A^{\prime}\smallsetminus B^{\prime}|=1$ since $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$. Applying Lemma˜5.5 to $(A^{\prime},B^{\prime})$ yields that $(A^{\prime},B^{\prime})$ is an atomic edge-cut.
˜(ii) $\Rightarrow$ ˜(i). From $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$ we get $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$. So no two edges in $S(A^{\prime},B^{\prime})$ share an end in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma˜5.5 applied to $(A^{\prime},B^{\prime})$. So suppose for a contradiction that there are two distinct edges $a_{1}b,a_{2}b$ in $S(A^{\prime},B^{\prime})$ with $a_{i}\in A^{\prime}\smallsetminus B^{\prime}$ and $b\in B^{\prime}\smallsetminus A^{\prime}$. Then $a_{i}\in A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $b\in B\cap A$. Hence $b$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, contradicting that $b$ was replaced with edges when taking the left-reduction. ∎
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ be a mixed-4-separation of $G$. Let $(C,D)$ be a 4-separation of $G$ with right-reduction $(C^{\prime},D^{\prime})$. Assume $A\subseteq C$. Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
@@
We obtain $A\subseteq C^{\prime}$ from $A\subseteq C=C^{\prime}$. Let $X$ consist of all endvertices in $A$ of edges in $S(A,B)$. Then $N(B\smallsetminus A)=(A\cap B)\cup X$. From $A\subseteq C$ we obtain $D\smallsetminus C\subseteq B\smallsetminus A$. Hence $D\subseteq(B\smallsetminus A)\cup N(B\smallsetminus A)=B\cup X$. Recall that $D^{\prime}\subseteq D$. Let $v\in D^{\prime}$. It suffices to show that $v\notin X$. If $v\in X$, then $v$ has exactly one neighbour in $B\smallsetminus A$. Since $B\smallsetminus A\supseteq D\smallsetminus C$, the definition of right-reduction yields $v\notin D^{\prime}$, a contradiction. So $v\notin X$. ∎
5.2. Left-right-reductions and right-left-reductions
@@
Let $(A,B)$ be a mixed-separation+. The left-right-reduction of $(A,B)$ is obtained from $(A,B)$ by first taking the left-reduction and then the right-reduction. The right-left-reduction of $(A,B)$ is obtained from $(A,B)$ by first taking the right-reduction and then the left-reduction.
@@
Let $(A^{\prime},B^{\prime})$ be the left-right-reduction of a mixed-4-separation+ $(A,B)$ in a graph $G$. Then $A^{\prime}\smallsetminus B^{\prime}\supseteq A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.
@@
We perform the two reductions one after the other, applying ˜5.2 after each reduction. ∎
@@
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$. The left-right-reduction $(A^{\prime\prime},B^{\prime\prime})$ of $(A,B)$ is a tetra-separation if and only if the following three conditions are satisfied:
@@
Let $(A^{\prime},B^{\prime})$ denote the left-reduction of $(A,B)$.
@@
Backward implication. Assume ˜(i)˜(iii). By ˜(iii) and Lemma˜5.4 (applied to $(B,A)$ instead of $(A,B)$), no two edges in $S(A,B)$ share an endvertex in $B\smallsetminus A$. By ˜(i) and Lemma˜5.6, $(A^{\prime},B^{\prime})$ satisfies the matching-condition. By Lemma˜5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$. By ˜(i) and since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$. Hence $(A^{\prime},B^{\prime})$ has order four by Lemma˜5.3. We further have $|A^{\prime}\smallsetminus B^{\prime}|=|A\smallsetminus B|\geqslant 2$ by ˜5.9 and we have $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$ by ˜(ii).
@@
By Lemma˜5.6 (applied for the right-reduction), $(A^{\prime\prime},B^{\prime\prime})$ satisfies the matching-condition. By Lemma˜5.3 (applied for the right-reduction), every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $B^{\prime\prime}\smallsetminus A^{\prime\prime}$. Every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $A^{\prime\prime}\smallsetminus B^{\prime\prime}$ since $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ by ˜5.9 (applied for the right-reduction). Hence $(A^{\prime\prime},B^{\prime\prime})$ satisfies the degree-condition. Since $B^{\prime}\smallsetminus A^{\prime}$ is nonempty and $G$ is 4-connected, every vertex in $S(A^{\prime},B^{\prime})$ has a neighbour in $B^{\prime}\smallsetminus A^{\prime}$, so $(A^{\prime\prime},B^{\prime\prime})$ has order four by Lemma˜5.3 (applied for the right-reduction). Finally, $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by ˜5.9, so $(A^{\prime\prime},B^{\prime\prime})$ is proper.
@@
Forward implication. Suppose first that ˜(iii) fails. Then, by Lemma˜5.4 for $(B,A)$, there are two edges $ab_{1},ab_{2}$ in $S(A,B)$ with $a\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$. Then $a\in A^{\prime\prime}\smallsetminus B^{\prime\prime}$ and $b_{i}\in B^{\prime\prime}\smallsetminus A^{\prime\prime}$ by ˜5.9. Hence $(A^{\prime\prime},B^{\prime\prime})$ fails the matching-condition.
@@
Suppose next that ˜(iii) holds, but ˜(i) or ˜(ii) fails. By ˜(iii) and Lemma˜5.4 for $(B,A)$, no two edges in $S(A^{\prime},B^{\prime})$ share an endvertex in $B^{\prime}\smallsetminus A^{\prime}$. If ˜(i) fails, we have $|A\smallsetminus B|\leqslant 1$, so Lemma˜5.6 gives that $(A^{\prime},B^{\prime})$ violates the matching-condition, and hence $(A^{\prime\prime},B^{\prime\prime})$ also violates the matching-condition. Next, assume that ˜(i) holds, but ˜(ii) fails. As $|A\smallsetminus B|\geqslant 2$, Lemma˜5.6 gives that $(A^{\prime},B^{\prime})$ satisfies the matching-condition. The failure of ˜(ii) yields $|B^{\prime}\smallsetminus A^{\prime}|\leqslant 1$. Hence applying Lemma˜5.6 to $(B^{\prime},A^{\prime})$ (instead of $(A,B)$) yields that $(A^{\prime\prime},B^{\prime\prime})$ violates the matching-condition. ∎
Lemma 5.11.
Let $k\in\mathbb{N}$. Let $G$ be a quasi-$(k+1)$-connected graph with $\geqslant k+4$ vertices. Then $G$ has no mixed-$k$-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
Proof.
@@
Assume that $G$ has a mixed-$k$-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Let $X:=A\smallsetminus B$ and $Y:=B\smallsetminus A$. Let $F$ denote the set of edges in $S(A,B)$. We have
$k+4\leqslant|G|=|X|+|Y|+k-|F|$
@@
Both $X$ and $Y$ have size $\geqslant 2$ by assumption. Hence both terms |X|2|X|-2 and |Y|2|Y|-2 are $\geqslant 0$. We may therefore bipartition $F$ into classes $F_{X}$ and $F_{Y}$ of $\leqslant|X|-2$ and $\leqslant|Y|-2$ edges, respectively. We obtain $X^{\prime}$ from $X$ by adding for each edge in $F_{X}$ its endvertex in $A$ to $X$, and we obtain $Y^{\prime}$ similarly. Then $(X^{\prime},Y^{\prime})$ is a $k$-separation of $G$. We have $|X^{\prime}\smallsetminus Y^{\prime}|\geqslant|X|-|F_{X}|\geqslant 2$, and similarly $|Y^{\prime}\smallsetminus X^{\prime}|\geqslant 2$. Thus $G$ is not quasi-$(k+1)$-connected. ∎
Proof of Lemma˜2.3..
˜(i). If $G$ is not quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Hence the right-left-reduction of $(A,B)$ is a tetra-separation of $G$ by ˜5.10.
˜(ii). Let $(A,B)$ be a tetra-separation of $G$. Then $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$ by Lemma˜4.2. Hence $G$ is not quasi-5-connected by Lemma˜5.11. ∎